OpenStudy (anonymous):

The base of a triangle is fixed and the ratio of the lenghts of the the other two sides is constant . Find the locus of the vertex joining the other two sides.

OpenStudy (anonymous):

\[\sqrt{x^2+y^2}=k \sqrt{(x-a)^2+y^2}\]

ganeshie8 (ganeshie8):

looks perfect !!

ganeshie8 (ganeshie8):

square on both sides n try to simplify...

OpenStudy (anonymous):

yeahh checked with my calc :) but have trouble turning that into circle form..

ganeshie8 (ganeshie8):

got it... gimme sometime to post the solution

ganeshie8 (ganeshie8):